OpenStudy (anonymous):

determine the convergence or divergence of the sequence with the given nth term. If the sequence converges, find its sum. a(sub)n= (n-2)!/n!

OpenStudy (anonymous):

start with \[\frac{(n-2)!}{n!}=\frac{1}{n(n-1)}\] and you should be in good shape

OpenStudy (anonymous):

i did that and then did partial fractions but i don't know how the sum equals one

OpenStudy (anonymous):

first off i hope we are starting at 2 because it is not defined if you start at n = 1

OpenStudy (anonymous):

it doesn't say in the problem

OpenStudy (anonymous):

but one and zero would both make a zero in the denominator so we'd have to start at two

OpenStudy (accessdenied):

Err, nevermind that. :p

OpenStudy (anonymous):

sorry that was a stupid question i asked, since the numerator is (n-2)!

OpenStudy (anonymous):

haha it's telescoping, i just didn't see that before. thank you!

OpenStudy (anonymous):

yes it sure is! you can see that from the partial fractions right?

OpenStudy (anonymous):

yeah, i just missed that before because i didn't do enough of the sums out. how do we know that it converges though?

OpenStudy (anonymous):

that pretty much tells you right there. of course could also compare to \[\sum\frac{1}{n^2}\] which you know converges

OpenStudy (anonymous):

or just say what the partial sums are

OpenStudy (anonymous):

yeah because p is greater than one. can you help me with \[\sum_{n=2}^{\infty} 1/n \sqrt{n ^{2}-1}\]

OpenStudy (anonymous):

you mean find the sum?

OpenStudy (anonymous):

determine if it converges or diverges and state the test used

OpenStudy (anonymous):

it is a set up for the integral test i believe

OpenStudy (anonymous):

i thought so but then i wasn't sure how to take the integral

OpenStudy (anonymous):

\[u=\sqrt{x^2-1}, du=\frac{x}{2\sqrt{x^2-1}}\]

OpenStudy (anonymous):

ok that was wrong, sorry

OpenStudy (anonymous):

for some reason i can't see the equation before i post

OpenStudy (anonymous):

oh that's weird

OpenStudy (anonymous):

but that is the idea, use a u - sub with \[u=\sqrt{x^2-1}\]

OpenStudy (anonymous):

but if you have the square root in the u- sub then you have to do chain rule

OpenStudy (anonymous):

yeah i messed up

OpenStudy (anonymous):

damn typos

OpenStudy (anonymous):

haha it's fine, you're being a huge help! i have a quiz tomorrow..

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

wait lets try this, since i can't type

OpenStudy (anonymous):

here, click on show steps and it will give you the integral. u sub for sure

OpenStudy (anonymous):

crap i still haven't memorized those arctans etc

OpenStudy (anonymous):

do you mind if i ask you another?

OpenStudy (anonymous):

hmm maybe another method but it would amount to a trig sub and you would get arctan in any case

OpenStudy (anonymous):

you can ask, but it might be best just to post an new question. you will get lots of responses

OpenStudy (anonymous):

i only need one right response haha

OpenStudy (anonymous):

Find the sum of the convergent series \[\sum_{n=1}^{\infty} 1/n(n+3)\]